# Taming Diffusion Transformer for Real-Time Mobile Video Generation

### v1

@@ Abstract
Diffusion Transformers (DiT) have shown strong performance in video generation tasks, but their high computational cost makes them impractical for resource-constrained devices like smartphones, and real-time generation is even more challenging. In this work, we propose a series of novel optimizations to significantly accelerate video generation and enable real-time performance on mobile platforms. First, we employ a highly compressed variational autoencoder (VAE) to reduce the dimensionality of the input data without sacrificing visual quality. Second, we introduce a KD-guided, sensitivity-aware tri-level pruning strategy to shrink the model size to suit mobile platform while preserving critical performance characteristics. Third, we develop an adversarial step distillation technique tailored for DiT, which allows us to reduce the number of inference steps to four. Combined, these optimizations enable our model to achieve over 10 frames per second (FPS) generation on an iPhone 16 Pro Max, demonstrating the feasibility of real-time, high-quality video generation on mobile devices.
@@ 1 Introduction
The rapid advancement of generative models [25, 6, 33] has led to significant breakthroughs in video generation [3, 29, 59, 44, 53, 42, 18], with Diffusion Transformers emerging as one of the most effective architectures for producing temporally coherent and visually compelling video content. These models leverage the strengths of diffusion processes [33, 12, 25] for stepwise refinement and transformer-based attention for capturing long-range dependencies across frames, making them particularly suitable for generating complex, high-fidelity video sequences. As such, they have become a cornerstone in state-of-the-art video synthesis pipelines.
@@
| Wan2.1 | 1.3 | 83.33 | 0.2 | |
| LTX | 1.8 | 80.00 | 6.1 | |
| Ours-Server | 2.0 | 83.09 | 6.4 | |
| Ours-Mobile | 0.9 | 81.45 | 151.3 | 12.2 |
Despite their impressive generative capabilities, Diffusion Transformers suffer from substantial computational overhead, especially when applied to high-resolution video generation. Though delivering a great quality boost, the computation and memory consumption of the 3D full attention [44, 53, 18, 9] scale quadratically with respect to total tokens (). This limitation poses a critical challenge for deploying these models in real-time or interactive settings, particularly on mobile devices with limited processing power and energy budgets. Existing efforts to optimize diffusion-based models, such as step reduction [57, 21, 54], efficient backbone [49, 51, 58], are mainly focused on Unet-based denoisers, which are naturally less expressive. Very few work [9] investigates the efficiency of DiT, and often suffers from perceptual quality or temporal consistency loss. Furthermore, most current acceleration methods are designed for server-class hardware [9, 44] and do not translate well to edge devices.
@@
In this work, we present a comprehensive optimization pipeline tailored specifically to accelerate video diffusion transformers for mobile deployment. Our approach combines three key strategies.
@@
(A) High-Compression Video Variational Autoencoder (VAE). First, we investigate the compression rate of the VAE. Video VAE with a higher compression ratio can significantly reduce the number of tokens in the latent representation, thus speeding up DiT inference. However, VAEs with an aggressive compression ratio often suffer from a loss of reconstruction quality, which likely leads to a loss in diffusion model generation quality. The trade-off between compression ratio and diffusion quality remains underexplored for on-device video generation. In this work, we create a series of video VAEs with different compression ratios and compare the speed gain versus generation quality loss. We have several findings: (i) the reconstruction and diffusion generation quality corresponds well with the compression ratio. (ii) The speedup from the higher VAE compression ratio is significant. (iii) Though slightly degraded, we can still find a sweet point that balances speed and quality.
@@
(B) Efficient Mobile DiT. Second, we find that directly training a lightweight DiT designated for mobile is challenging. Instead, we start from a larger pre-trained supernet and propose a sensitivity aware tri-level pruning with KD-Guided framework that selectively removes less critical components of the model based on their contribution to both runtime and output quality. This pruning reduces the number of DiT blocks, feed-forward features, and attention heads. The final architecture has M parameters and can be easily deployed on a modern device such as an iPhone 16 Pro Max. Further, we improve the pruned model performance by aligning features of the pruned network and the supernet through knowledge distillation.
@@
(C) Adversarial Step Distillation. Finally, we introduce a novel adversarial distillation method for DiTs. It allows us to achieve comparable quality to full-step diffusion model using only a small number of inference steps, dramatically reducing computational cost. Prior adversarial distillation methods for video diffusion models mainly focus on Unet backbones [51, 57] and less challenging image-to-video tasks [3]. We find that they cannot be readily applied to diffusion transformers. We design a new discriminator architecture where we inherit the first blocks from the generator and freeze them as the time-conditioned feature parser, and add both full 3D attention and cross attention as learnable discriminator layers. Our design enables step generation without classifier-free guidance (CFG), which is faster compared with a typical -step recipe with CFG.
@@
With these optimizations, our model can generate high-quality video at over frames per second (FPS) on an iPhone 16 Pro Max using only four denoising steps. Extensive experiments demonstrate that our method maintains strong visual fidelity and temporal consistency, closely matching the outputs of full-resolution, unpruned models. For the first time, our work advances the state-of-the-art for on-device efficient video generation by making real-time diffusion-based video synthesis feasible on consumer-grade mobile hardware. Our contributions can be summarized as follows,
@@
To obtain an efficient DiT backbone, we find that training a smaller network from scratch yields inferior generation results. We instead start from a large pre-trained super network and perform distillation-guided, sensitivity-aware pruning. We deliver a compact network with optimized depth and width configuration.
@@ 2 Related Work
Video Diffusion Model
There has been an exponential growth in the literature for video generation models [44, 29, 24, 53, 18, 42, 19, 26]. Most of the recent efforts have been towards developing large video diffusion models, which learn to iteratively denoise a pure Gaussian noise into a semantically meaningful and content-rich video sample conditioned on user inputs such as text-prompt or image. These include pixel space [28, 11] and latent space [44, 53] video models. While these models [29, 59, 28, 32, 9, 53] are capable of generating realistic and visually appealing videos, their resource requirements render them unsuitable for on-device deployment.
There have been very little efforts towards on-device video generation [49, 17]. Wan2.1 [44] family includes a B T2V model, but due to low VAE compression ratio, it yields large number of latent tokens prohibiting the deployment on-device. LTX-Video [10] model uses a high-compression VAE, resulting in compressed latents. Although LTX-Video generates videos in real-time and is considerably faster than Wan2.1, their model size (B) is still too restrictive to be deployed on-device. SnapGen-V [49] utilizes a tiny UNet denoiser for diffusion modeling, but the generated videos have poor quality. Mobile Video Diffusion [51] prunes the Stable Video Diffusion [3] model and reduces the number of channels and UNet blocks. On-device Sora [17] enables low resolution video generation on iPhones by enabling token merging in temporal dimension and concurrent model block loading to address the limited working memory. In this work, we propose an efficient mobile DiT architecture and utilize a high-compression VAE to design an on-device real-time video generation model which generates high-quality high-resolution videos on mobile devices such as iPhone 16 Pro.
Step Distillation.
Since diffusion models [6, 31, 13] iteratively denoise the gaussian noise, they require substantial number of steps, each requiring a forward pass through a large neural network, to generate a video. Thus, reducing the number of inference steps directly cuts down the video generation latency. There have many works [55, 54, 52, 47, 40, 39, 57, 49] that aims to solve this problem of step distillation in diffusion models. Most of these works have been for the text-to-image models [55, 54, 52, 47, 16, 27, 5].
Early works [34, 22] proposed progressive distillation wherein a student model learns to predict the output of two inference steps of a pretrained teacher. This procedure when repeated yields a few step diffusion model, but this requires progressively refining the inference steps via a new training procedure, resulting in computationally expensive step distillation approach. Consistency Models [40, 39] refine the distillation objective to enforce a consistency property where each point in trajectory maps to the clean data. Later works [50, 38, 35] further incorporate adversarial loss to distill a single-step student, and enhance multi-step results as well. Recently proposed Shortcut models [8] introduce flow-matching loss with self-consistency loss while modifying the denoiser to be aware of the inference step.
In the context of video models, there have been works such as SF-V [57] and SnapGen-V [49] that enables few step video generation through step distillation. SF-V [57] enables one-step video generation for the SVD [3] model by adding latent adversarial training. Inspired by SF-V, SnapGen-V [49] proposes a few step video generation model for an efficient UNet denoiser by developing a unified spatio-temporal discriminator design in the adversarial training.
@@ 3 Preliminaries
We recap the basics of video diffusion models and define our annotations. Following popular practices of latent diffusion [59], we employ a video autoencoder to encode video data into a compressed latent space , where is the number of temporal frames, and are the spatial resolutions, and is the latent channels. The VAE compression ratio is thus , e.g., [53, 18, 44] and [2] VAEs. The objective of the DiT generator is to generate under certain guidance (i.e., text prompt).
We employ Rectified Flow [47] to train our latent DiT model. According to the flow-matching-based diffusion process, given a clean video latent , the intermediate noisy state at a timestep is:
@@ 4.1 Scaling Latent Compression Ratio
DiT demonstrates superior generation capabilities when attending on full token length () [53], however, it is also notorious for its quadratic computational cost. The key idea of the latent diffusion model is to construct a compressed latent space and reduce the generation cost. As a result, a straightforward idea to accelerate DiT is to further increase the VAE compression ratio. State-of-the-art models (CogVideoX [53],Hunyuan[18],Wan [44]) employ a VAE combined with a patchify module, which comprises a total compression rate, while the recent OpenSora-2 [59] adopts a VAE, and LTX [9] adopts an VAE to reduce the dimensionality of the latent features input to the DiT and results in faster generation speed. However, there has been limited research on how the VAE compression ratio affects the quality and speed of video generation. Upon aggressive compression, it becomes more challenging for the VAE decoder to fully reconstruct the details, which may result in quality loss. In this work, we perform a comprehensive study on the scaling of the VAE compression ratio. We follow [9, 48] and construct video VAEs with various compression ratios from to . We build the VAE with 3D convolutions to better handle video modality, and use a fixed latent channel number for all variants. We train the same DiT network under each latent space and benchmark the generation speed and quality. Detailed results and discussions can be found in Sec. 5.3.
@@ 4.2 Efficient DiT Architecture via KD-Guided Tri-Level Pruning
Despite operating in a highly compressed latent space, the size of the Diffusion Transformer (DiT) remains a critical factor in edge generation scenarios [49], where mobile devices are constrained by limited memory, power, and computational resources. Training a compact DiT that still achieves high-quality generation is a non-trivial challenge. First, DiT models generally exhibit strong generation capabilities only when scaled to a sufficiently large capacity. Moreover, designing an effective small-scale DiT is difficult due to the high-dimensional design space—including network depth (number of transformer blocks), width (channel size), and attention head count.
A promising alternative is to begin with a well-trained large model and prune it to meet resource constraints. Prior work such as TinyFusion [7] explored this approach via block-wise pruning using a learnable layer mask, effectively constructing a shallower DiT. However, as the properties demonstrated in Figure 3, it is still of great value to design a fine-grained pruning method that offers deeper insight into which parameters are critical or redundant, thereby enabling a better trade-off between efficiency and generation quality.
@@ 4.2.1 Tri-level Pruning
As exhibited in Figure 3, the transformer block pruning is a simple yet coarse approach, we consider it a low-granularity pruning scheme. To enable finer granularity and better address redundancies, we propose a tri-level pruning scheme that incorporates block pruning and further introduces fine-grained pruning techniques, including head pruning for the multi-head attention mechanism and channel pruning for the linear layer. Notably, the pruned model can be converted into a dense and compact form, enabling execution on mobile devices without requiring additional compilation or specialized hardware support.
@@
We employ a set of learnable binary masks to implement the tri-level pruning scheme. Each binary mask encodes the importance of its corresponding granularity, i.e. block, attention-head, or linear dimension. A mask value of 0 indicates that the corresponding unit should be pruned, while a value of 1 denotes that it should be preserved. Specifically, block pruning can be formulated as shown in Eq. 3, where indicate the input and output features of the DiT block, is the binary mask associated with that block, and denotes the set of binary masks for all DiT blocks. When , the block is active; otherwise, its output is bypassed through a residual connection.
@@
The other two pruning schemes can be expressed using a unified formulation, since pruning attention heads is equivalent to removing specific output features before the multi-head attention operation for each token. By integrating the pruning mechanism into the linear layer, the operation can be formulated as shown in Eq. 4:
@@
where denote the input and output features of the linear layer, and is a binary mask with -dimension corresponding to the output channels. For each , a value of zeros out the corresponding output channel at dimension for layer ; otherwise the channel remains active.
@@
The proposed tri-level pruning scheme begins by generating a candidate mask set for each pruning target, as illustrated in Figure 2. These candidate masks are selected based on the desired number of active components, which are constrained by the memory limitation of the target device. Notably, exhaustively exploring all pruning combinations results in an extremely large search space, making the optimization problem intractable (e.g., pruning 6 out of 32 attention heads results in 906,192 possible configurations). To mitigate this, we adopt a group-wise masking mechanism that partitions overall search space into smaller subspaces, allowing pruning to be performed efficiently within each subspace. Once the candidate masks are generated, we further optimize them to identify the optimal configuration that minimizes the information loss caused by pruning.
Here, we specify the details of our tri-level pruning scheme for constructing an efficient Diffusion Transformer architecture tailored to the iPhone 16 Pro Max. Due to the memory limitation of the device, the total number of parameters must remain under 1 billion. Based on the sensitivity analysis in Figure 3, which indicates that FFN contributes more significantly to performance than attention heads, we prioritize pruning attention heads more aggressively. Starting from a 2B parameter base model with 28 DiT blocks, 32 attention-heads, and FFN dimension of 8192, our final efficient model archives 915M parameters by pruning 8 blocks, 12 attention heads, and reducing the FFN dimension by 25%. Detailed results of the pruned model can be found in Sec. 5.

### v2

# Taming Diffusion Transformer for Efficient Mobile Video Generation in Seconds

@@ Abstract
Diffusion Transformers (DiT) have shown strong performance in video generation tasks, but their high computational cost makes them impractical for resource-constrained devices like smartphones, and practical on-device generation is even more challenging. In this work, we propose a series of novel optimizations to significantly accelerate video generation and enable practical deployment on mobile platforms. First, we employ a highly compressed variational autoencoder (VAE) to reduce the dimensionality of the input data without sacrificing visual quality. Second, we introduce a KD-guided, sensitivity-aware tri-level pruning strategy to shrink the model size to suit mobile platforms while preserving critical performance characteristics. Third, we develop an adversarial step distillation technique tailored for DiT, which allows us to reduce the number of inference steps to four. Combined, these optimizations enable our model to achieve approximately 15 frames per second (FPS) generation speed on an iPhone 16 Pro Max, demonstrating the feasibility of efficient, high-quality video generation on mobile devices.
@@ 1 Introduction
The rapid advancement of generative models (Liu et al., 2022; Esser et al., 2024; Rombach et al., 2022) has led to significant breakthroughs in video generation (Blattmann et al., 2023; OpenAI, 2023; Zheng et al., 2024; Wan et al., 2025; Yang et al., 2024b; Team, 2024; Kong et al., 2024), with Diffusion Transformers emerging as one of the most effective architectures for producing temporally coherent and visually compelling video content. These models leverage the strengths of diffusion processes (Rombach et al., 2022; Ho et al., 2020; Liu et al., 2022) for stepwise refinement and transformer-based attention for capturing long-range dependencies across frames, making them particularly suitable for generating complex, high-fidelity video sequences. As such, they have become a cornerstone in state-of-the-art video synthesis pipelines.
Despite their impressive generative capabilities, Diffusion Transformers suffer from substantial computational overhead, especially when applied to high-resolution video generation. While this brings significant quality improvements, the computation and memory consumption of the 3D full attention (Yang et al., 2024b; Kong et al., 2024) scale quadratically with respect to total tokens (). This limitation poses a critical challenge for deploying these models in interactive settings, particularly on mobile devices with limited processing power and energy budgets. Existing efforts to optimize diffusion-based models, such as step reduction (Zhang et al., 2024; Li et al., 2023; Yin et al., 2024a), efficient backbone (Wu et al., 2025b; Yahia et al., 2024; Zhao et al., 2024), are mainly focused on UNet-based denoisers, which are naturally less expressive. Very few work (HaCohen et al., 2024b) investigates the efficiency of DiT, and often suffers from perceptual quality or temporal consistency loss. Furthermore, most current acceleration methods are designed for desktop (HaCohen et al., 2024b; Wan et al., 2025), and do not translate well to edge devices.
@@
| Wan2.1 | 1.3 | 83.33 | 0.2 | ✗ |
| LTX | 1.8 | 80.00 | 6.1 | ✗ |
| Ours-Server | 2.0 | 83.09 | 6.4 | ✗ |
| Ours-Mobile | 0.9 | 81.45 | 151.3 | 15 |
@@
In this work, we present a comprehensive optimization pipeline tailored specifically to accelerate video diffusion transformers for mobile deployment. Our approach combines four key strategies.
@@
(A) High-Compression Video Variational Autoencoder (VAE). First, we investigate the compression rate of the VAE. High compression video VAE can significantly reduce the number of tokens in the latent representation, thus speeding up DiT inference. However, VAEs with an aggressive compression ratio often suffer from a loss of reconstruction quality, which likely leads to a loss in diffusion model generation quality. The trade-off between compression ratio and diffusion quality remains underexplored for on-device video generation. In this work, we create a series of video VAEs with different compression ratios and compare the speed gain versus generation quality loss. We have several findings: (i) The reconstruction and diffusion generation quality corresponds well with the compression ratio. (ii) The speedup from the higher VAE compression ratio is significant. (iii) Though slightly degraded, we can still find a sweet point that balances speed and quality.
@@
(B) Efficient Mobile DiT. Second, we find that directly training a lightweight DiT designated for mobile is challenging. Instead, we start from a larger pre-trained supernet and propose a sensitivity-aware tri-level pruning with a KD-Guided framework that selectively removes less critical components of the model based on their contribution to both runtime and output quality. This pruning reduces the number of DiT blocks, feed-forward features, and attention heads. The final architecture has M parameters and can be easily deployed on a modern device such as an iPhone 16 Pro Max. Further, we improve the pruned model performance by aligning features of the pruned network and the supernet through knowledge distillation.
@@
(C) Adversarial Step Distillation. Third, we design a new discriminator head tailored for adversarial step-distillation on DiTs that achieves full-step quality with only a few sampling steps. Prior adversarial distillation methods for video diffusion models mainly focus on UNet backbones (Yahia et al., 2024; Zhang et al., 2024) and less challenging image-to-video tasks (Blattmann et al., 2023) and do not transfer directly to DiTs. Our discriminator design inherits the first frozen pretrained generator blocks as a time-conditioned feature parser, and adds learnable with self-attention and cross-attention to fully capture conditions. This design enables four-step generation without classifier-free guidance (CFG), yielding faster inference than a typical 40-step CFG recipe.
(D) Operator Optimization for Efficient Inference. Finally, we identify a memory bottleneck in the linear layer of DiTs (i.e. feed-forward network), where the limited bandwidth prevents operators from approaching their theoretical speed on device. To address this, we introduce a tiled GEMM strategy that alleviates the memory bottleneck without requiring kernel-level modifications. The design achieves over 50% speedup on targeted linear layers and 10% acceleration for DiT inference.
@@
With these optimizations, our model can generate high-quality video at over frames per second (FPS) on an iPhone 16 Pro Max using only four denoising steps. Extensive experiments demonstrate that our method maintains strong visual fidelity and temporal consistency, closely matching the outputs of full-resolution, unpruned models. For the first time, our work advances the state-of-the-art for on-device efficient video generation by making practical diffusion-based video synthesis feasible on consumer-grade mobile hardware. Our contributions can be summarized as follows,
@@
To obtain an efficient DiT backbone, training a smaller network from scratch gives inferior results. Instead, we start from a large pre-trained super network and apply distillation-guided, sensitivity-aware pruning, yielding a compact network with optimized depth and width.
@@
We identify the memory bottleneck in the feed-forward layers of DiTs on device performance and introduce a tiled GEMM strategy that alleviates this issue, enabling more efficient and hardware friendly on-device inference.
@@ 2 Related Work
Video Diffusion Models. Recent years have seen rapid progress in video generation models (Wan et al., 2025; OpenAI, 2023; Lin et al., 2024; Yang et al., 2024b; Kong et al., 2024; Team, 2024; Kuaishou, 2024; Ma et al., 2025). Most advances focus on large diffusion models that iteratively denoise Gaussian noise into realistic videos, conditioned on text or images. These approaches include pixel-space models (Menapace et al., 2024; Ho et al., 2022) and latent-space models (Wan et al., 2025; Yang et al., 2024b). While such systems (OpenAI, 2023; Zheng et al., 2024; Menapace et al., 2024; Polyak et al., 2024; HaCohen et al., 2024b; Yang et al., 2024b) generate high-quality videos, their resource demands make them unsuitable for on-device use.
On-Device Models. In contrast, only limited work targets on-device video generation (Wu et al., 2025b; Kim et al., 2025). The Wan2.1 family (Wan et al., 2025) includes a B T2V model, but its low VAE compression yields too many latent tokens for deployment. LTX-Video (HaCohen et al., 2024a) applies a high-compression VAE and runs in real time on GPUs, yet its B parameters remain prohibitive for mobile devices. SnapGen-V (Wu et al., 2025b) adopts a lightweight UNet but sacrifices visual fidelity. Mobile Video Diffusion (Yahia et al., 2024) reduces Stable Video Diffusion (Blattmann et al., 2023) by pruning channels and blocks. On-device Sora (Kim et al., 2025) achieves low-resolution video generation on iPhones via temporal token merging and concurrent block loading to handle memory limits.
Step Distillation. Diffusion models (Esser et al., 2024; Podell et al., 2023; Hoogeboom et al., 2023) require many denoising steps, each involving a full network pass, which creates latency. Reducing the number of steps directly improves efficiency. Numerous methods address this in text-to-image tasks (Yin et al., 2024b; a; Yang et al., 2024a; Wang et al., 2024b; Kim et al., 2024; Mei et al., 2024; Dao et al., 2025), representative work including progressive distillation (Salimans & Ho, 2022; Li et al., 2023), consistency models (Song et al., 2023; Song & Dhariwal, 2023), adversarial training (Xu et al., 2024; Sauer et al., 2023b; 2024), shortcut models (Frans et al., 2024), and mean flow (Geng et al., 2025). For video, Zhang et al. (2024); Wu et al. (2025b) achieves few-step generation Blattmann et al. (2023) with adversarial training specially-designed spatio-temporal discriminator.
@@ 3 Preliminaries
Following popular practices of latent diffusion (Zheng et al., 2024), we employ a video autoencoder to encode video data into a compressed latent space , where is the number of temporal frames, and are the spatial resolutions, and is the latent channels. The VAE compression ratio is thus , e.g., (Yang et al., 2024b; Kong et al., 2024; Wan et al., 2025) and (Agarwal et al., 2025) VAEs. The objective of the DiT generator is to generate under certain guidance (i.e., text prompt).
We employ Rectified Flow (Wang et al., 2024b) to train our latent DiT model. According to the flow-matching-based diffusion process, given a clean video latent , the intermediate noisy state at a timestep is:
@@ 4 Method
We optimize Diffusion Transformer (DiTs) for efficient on-device video generation from four perspectives: (a) High compress VAE:we employ a high-compression autoencoder, as the computational complexity of transformer scales quadratically with token length. Reducing the number of token decreases computation while enabling video generation at higher resolution and longer duration. (b) Efficient DiT architecture:we design an efficient DiT using a KD-Guided Tri-Level pruning method. The method balances model fidelity to the baseline with the hardware constraint of the target device. (c) Step-distillation:we adopt adversarial fine-tuning for step distillation with a new discriminator head design, reducing the number of sampling steps and achieving up to acceleration during inference. (d) Operator Optimization:we identify the memory bottlenecks in DiT feed-forward layers and introduce a tiled GEMM strategies, alleviating bandwidth limitations and enabling efficient on-device inference without requiring kernel/compiler modification.
@@ 4.1 Scaling Latent Compression Ratio
DiT demonstrates superior generation capabilities when attending on full token length (), however, it is also notorious for its quadratic computational cost. The key idea of the latent diffusion model is to construct a compressed latent space and reduce the generation cost. As a result, a straightforward idea to accelerate DiT is to further increase the VAE compression ratio. State-of-the-art models (CogVideoX (Yang et al., 2024b),Hunyuan(Kong et al., 2024),Wan (Wan et al., 2025)) employ a VAE combined with a patchify module, which comprises a total compression rate, while the recent OpenSora-2 (Zheng et al., 2024) adopts a VAE, and LTX (HaCohen et al., 2024b) adopts an VAE to reduce the dimensionality of the latent features input to the DiT and results in faster generation speed. However, there has been limited research on how the VAE compression ratio affects the quality and speed of video generation. Upon aggressive compression, it becomes more challenging for the VAE decoder to fully reconstruct the details, which may result in quality loss. In this work, we perform a comprehensive study on the scaling of the VAE compression ratio. We follow HaCohen et al. (2024b); Wu et al. (2025a) and construct video VAEs with various compression ratios from to . We build the VAE with 3D convolutions to better handle video modality, and use a fixed latent channel number for all variants. We train the same DiT network under each latent space and benchmark the generation speed and quality. Results and discussions are in Secs.˜5.3 and G.
@@ 4.2 Efficient DiT Architecture via KD-Guided Tri-Level Pruning
Despite operating in a highly compressed latent space, the size of the Diffusion Transformer (DiT) remains a critical factor in edge generation scenarios (Wu et al., 2025b), where mobile devices are constrained by limited memory, power, and computational resources. Training a compact DiT that still achieves high-quality generation is a non-trivial challenge. First, DiT models generally exhibit strong generation capabilities only when scaled to a sufficiently large capacity. Moreover, designing an effective small-scale DiT is difficult due to the high-dimensional design space—including network depth (number of transformer blocks), width (channel size), and attention head count.
A promising alternative is to begin with a well-trained large model and prune it to meet resource constraints. Prior work such as TinyFusion (Fang et al., 2024) explored this approach via block-wise pruning using a learnable layer mask, effectively constructing a shallower DiT. However, as the properties demonstrated in Figure˜3, it is still of great value to design a fine-grained pruning method that offers deeper insight into which parameters are critical or redundant, thereby enabling a better trade-off between efficiency and generation quality.
@@ 4.2.1 Tri-level Pruning
As exhibited in Figure˜3, the transformer block pruning is a simple yet coarse approach, we consider it a low-granularity pruning. To enable finer granularity and better address redundancies, we propose a tri-level pruning scheme that incorporates block pruning and further introduces fine-grained pruning techniques, including head pruning for the multi-head attention mechanism and channel pruning for the linear layer. Notably, the pruned model can be converted into a dense and compact form, enabling execution on mobile devices without requiring additional compilation or specialized hardware support.
@@
We employ a set of learnable binary masks to implement the tri-level pruning scheme. Each binary mask encodes the importance of its corresponding granularity, i.e. block, attention-head, or linear dimension. A mask value of 0 indicates that the corresponding unit should be pruned, while a value of 1 denotes that it should be preserved. Specifically, block pruning can be formulated as shown in Eq.˜3, where indicate the input and output features of the DiT block, is the binary mask associated with that block, and denotes the set of binary masks for all DiT blocks. When , the block is active; otherwise, its output is bypassed through a residual connection.
@@
The other two pruning schemes can be expressed using a unified formulation, since pruning attention heads is equivalent to removing specific output features before the multi-head attention operation for each token. By integrating the pruning mechanism into the linear layer, the operation can be formulated as shown in Eq.˜4:
@@
where denote the output and input features of the linear layer, and is a binary mask with -dimension corresponding to the output channels. For each , a value of zeros out the corresponding output channel at dimension for layer ; otherwise the channel remains active.
@@
The proposed tri-level pruning scheme begins by generating a candidate mask set for each pruning target, as illustrated in Figure˜2. These candidate masks are selected based on the desired number of active components, which are constrained by the memory limitation of the target device. Notably, exhaustively exploring all pruning combinations results in an extremely large search space, making the optimization problem intractable (e.g., pruning 6 out of 32 attention heads results in 906,192 possible configurations). To mitigate this, we adopt a group-wise masking mechanism that partitions overall search space into smaller subspaces, allowing pruning to be performed efficiently within each subspace. Once the candidate masks are generated, we further optimize them to identify the optimal configuration that minimizes the information loss caused by pruning.